2. A study shows that length of the careers of professional basketball players are nearly normally distributed, with a mean of 6.1 years and a standard deviation of 1.8 years.

1
Expert's answer
2021-11-03T08:50:48-0400

Question is incomplete. We can add last line as ''What percent of professional basketball players have a career of more than 9 years?

Solution:

"\text{Given: Mean }(\mu) = 6.1\ \text{ year}
\\\text{Standard deviation}\ (\sigma)=1.8
\\\begin{aligned}
\\P(x>9) &=1-P(x<9) \\
\\&=1-p\left(\frac{x- \mu}{\sigma}<\frac{9-6.1}{1.8}\right) \\
\\&=1-P\left(z<\frac{2.9}{1.8}\right) \\
\\&=1-P(z<1.61) \\
\\&=1-0.9463 \\
\\&=0.0537 \\
\\&=5.37 \%
\end{aligned}"

"\\\therefore 5.37 \%" is the percent of professional basketball player have a career of more than 9 years.